question 45

Etling Inc.'s dividend is expected to grow at 7% for the next two years and then at 3% forever. If the current dividend is $3 and the required return is 16%, what is the price of the stock?

Select one:

a. $25.10

b. $26.15

c. $27.58

d. $25.54

e. $29.45

question 39

This morning, Alicia bought a ten-year 6% coupon bond that pays interest semi-annually. She paid $994 for a $1,000 bond. If the market interest rate on this type of bond declines to 5.5% tonight, how much will Alicia receive for her first interest payment?

Select one:

a. $60.00

b. $35.00

c. $69.58

d. $30.00

e. $65.00
Answer #1

D1 = 3 * 1.07

D2 = 3 * 1.07 * 1.07

D3 = 3 * 1.07 * 1.07 * 1.03

P2 = D3/ (Required rate - Growth rate)

P2 = 3 * 1.07 * 1.07 * 1.03/ (16% - 3%)

P2 = 27.21

P0 = 3 * 1.07/ (1.16) + 3 * 1.07 * 1.07/ (1.16)2 + 27.21/ (1.16)2

P0 = 2.77 + 2.55 + 20.22

P0 = 25.54

Option D is correct.

Question 39:

As coupon payment does not depend on the market rate, it will not change with increase or decrease of market interest rate.

Its coupon payment will remain constant, which will be:

PMT = 6% * 1,000/ 2

Above, we divided by 2 to calculate semi annual payment

PMT = 30

Option d is correct